undeadknight26 (undeadknight26):

What is the equation in standard form of the line which passes through (4, −2) and has a slope of −3?

undeadknight26 (undeadknight26):

10 = -3x + x?

OpenStudy (eric_d):

You may use y-y1=m(x-x1)

OpenStudy (eric_d):

Just plug in the given value.. m- slope/gradient

undeadknight26 (undeadknight26):

y - y1 = m(x - x1) y - -2 = -3(x - 4) @eric_d now just solve?

OpenStudy (eric_d):

yes

undeadknight26 (undeadknight26):

Not sure where to actually go from here though sadly o.o

OpenStudy (eric_d):

Then arrange it into this form.. y=mX+b

OpenStudy (eric_d):

First, you'll need to expand that

OpenStudy (eric_d):



OpenStudy (eric_d):

Then, arrange it into this form.. y=mX+b

OpenStudy (eric_d):

b= y-intercept

OpenStudy (eric_d):

Do you get me @undeadknight26

undeadknight26 (undeadknight26):

Hmm i think so...

OpenStudy (eric_d):

Okay, good !

undeadknight26 (undeadknight26):

y + 2 = -3x + 12? y + -3x = 2 + 12 or no?

OpenStudy (eric_d):

Nope

OpenStudy (eric_d):

You'll need to arange that into y=mX+c

OpenStudy (eric_d):

or y=mX+b It's the same b=c=y-intercept

OpenStudy (eric_d):